# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v4

@@ 2.3 Hard negative mining
In recent years, deep learning-based object detection models have also adopted this strategy, reporting significant performance improvements by relearning misdetected bounding box regions, including in the medical field (Li et al., 2019; Tang et al., 2019). However, such object detection models typically cannot utilize images that do not contain any detection targets (such as background images or healthy control images) for training, since the error function to be minimized is defined based on the position and size of the bounding box for each detection target. Existing hard sample mining methods have been applied to data with at least one annotation box of the target class and have used overdetected or undetected regions or image patches as hard samples. Our HBBT is a form of hard sample mining or hard negative mining, which retrains regions of false positives. However, it differs significantly in that it allows the use of healthy control cases, which could not be used in traditional object detection models, for training. This is a particularly important contribution in the medical field, where the cost of acquiring positive samples is high.
3 Practical gastric cancer screening system
This study proposed a practical gastric cancer screening system using gastric X-ray images. The system highlights suspected cancerous sites with bounding boxes, allowing physicians to confirm the validity of the estimations. The two technical innovations in this study significantly improved system performance and achieved practical accuracy. The overall structure of the proposed system is shown in Figure 1. This study was approved by the Institutional Review Board of Tokai University Hospital (No. 20R-033, approval date: June 12, 2020). The main technical contributions were (1) R-sGAIA, which generates a diverse set of images that effectively emphasize the gastric fold region for data augmentation, and (2) HBBT, which reduces false positives by assigning “hard-sample” classes to bounding box regions where false positives are detected and iteratively learning from them. The object detection model used is EfficientDet (Tan et al., 2020), which incorporates EfficientNet (Tan and Le, 2019), a CNN known for its excellent performance, as its backbone, although the choice of model was arbitrary. The input images were 2,048 $\times$ 2,048 pixels, and the output consisted of a bounding box surrounding the presumed cancerous region in the image along with its confidence score.
Figure 1: Overview of the proposed gastric cancer screening system.
@@ 3.1 R-sGAIA
The proposed R-sGAIA is a refined version of our previous method, sGAIA (Okamoto et al., 2019). Both R-sGAIA and sGAIA are probabilistic online data augmentation techniques designed to generate various images that highlight gastric folds in X-ray images, using medical knowledge to detect inflammation on the gastric mucosal surface. The original sGAIA improved gastric cancer detection by 6.9% in terms of F1-score (recall 92.3%, precision 32.4%) when applied to a Faster R-CNN model with ResNet-101 as the backbone, validating its effectiveness for practical screening. However, in sGAIA, the emphasis probability of each pixel $p(x,y)$, crucial for effective enhancement, was subjectively determined as a discrete value based on preliminary experiments. In the proposed R-sGAIA, $p(x,y)$ is instead modeled as a flexible continuous function, determined by only two hyperparameters. A schematic of the proposed R-sGAIA process is shown in Figure 2. Similar to sGAIA, R-sGAIA consists of four steps, with step 2 being the only difference.
@@
    (Step 1) Calculation of edge strength $E(x,y)$
    First, a contrast-enhanced image $I_{e}(x,y)$ is generated from the original gastric X-ray image $I(x,y)$ using histogram equalization. Then, its gradient and high-frequency components, $\nabla I_{e}(x,y)$, and $H_{e}(x,y)$, are obtained using general image processing techniques such as Canny edge detection filter and Butterworth high-pass filter, respectively. Each of them is normalized to [0, 1], and the normalized edge strength $E(x,y)$ is calculated as follows:
@@
    (Step 2) Calculation of the probability of a gastric fold region $p(x,y)$
@@
    The parameters $\gamma$ and $\theta$ are hyperparameters, respectively, defining the slope of the sigmoid function and adjusting the probability of running the augmentation of $E(x,y)$ to be 50%. They are determined based on the diagnostic performance of the validation data, as described below. Figure 3 compares the enhancement selection probability $p(e=E(x,y))$ based on the feature intensity $e$ for R-sGAIA and sGAIA. Note that, in sGAIA, the probability $p(e)$ is subjectively determined based on preliminary experiments, as described above.
@@
    (Step 3) Determination of gastric fold edge region $G(x,y)$
@@
    (Step 4) Generation of enhanced gastric fold images $A(x,y)$
@@
Here, $\alpha$ and $\beta$ are augmentation parameters. The value of $\alpha$ ranges from 0.9 to 1.0, and $\beta$ ranges from --15 to --5, with both values chosen randomly for each processing. Based on this, the proposed R-sGAIA probabilistically enhances the location and intensity of the gastric fold region each time, generating a different enhanced gastric image for each trial. This makes it highly compatible with online data augmentation in deep neural network training. Another significant advantage of R-sGAIA is that it does not require any additional runtime. We applied this R-sGAIA for online augmentation in training a machine learning model (EfficientDet-D7) used for cancer region detection.
@@ 3.2 HBBT
The diagnostic support system for gastric X-ray images is a one-class detection problem focused solely on positive (i.e., malignant) regions. However, our task involves many regions with similar image features, making them difficult to distinguish, even for experts. With a large number of training images, recent deep learning-based object detection algorithms can achieve the desired detection performance for this task. However, this is challenging when labeled supervisory data is limited. Figure 4 shows typical false positives identified by experts in experiments using a vanilla detection model (i.e., EfficientDet-D7 without the proposed HBBT). The results indicate that most of these false positives fall into the five categories shown in examples (b)–(f) in Figure 4. However, it is difficult to clearly define and control these features in advance. In a fine-grained task like ours, it is crucial to reduce such hard negatives that exhibit high image similarity but are not the target. Note that our system detects a single class of candidate gastric cancer regions, with each detection box labeled as ”malignant.”
One promising approach to address this issue is hard sample mining, specifically hard negative mining, which aims to reduce overdetection and has been extensively studied, as mentioned earlier. However, training hard negative mining in object detection models typically focuses on reducing false positives (hard negatives) outside the target region in the training image. The data available for training are constrained by the object detection model, meaning that only images with some object detection annotation label can be used, unless a framework like semi-supervised learning is employed.
As mentioned earlier, a significant area for improvement in object detection models is the active use of negative samples, i.e., healthy controls. Conventional deep learning-based object detection algorithms typically only use images containing the object to be detected, excluding control cases without the object from training. This is a critical limitation. To address this, we propose HBBT, a simple yet effective method for suppressing false positives by incorporating the concept of hard negative mining. Training negative samples, which are more prone to false positives among the training data, is generally effective in constructing detectors with fewer false positives. In HBBT, as shown in Figure 1, the false-positive bounding boxes detected by the model are called hard boundary boxes. These are assigned hard-sample labels, converting the task into a two-class detection problem. The newly added hard-sample class label is treated as a negative example, and training is repeated until an optimal F1-score is achieved on the validation data. Importantly, this approach allows the use of not only positive data (e.g., images with cancerous regions) but also negative (i.e., healthy control) data, which are generally available in large quantities, for training by detecting and utilizing hard boundary boxes.
Figure 4: Positive example and five typical false positives provided by experts (as shown in red box).
@@ 4.1 Datasets and evaluation criteria
The image set used in this study consisted of 4,724 gastric X-ray images (1,117 images with 1,504 lesion annotations by radiologists and 3,607 images without lesions) from 145 patients in clinical settings, provided by the Tokai University School of Medicine, Japan. A total of 116 images from five randomly selected patients (49 of them with 49 annotated boxes) were used as validation data to determine the hyperparameters and to decide when to terminate HBBT. The remaining 4,608 images from 140 patients were used to train and evaluate the proposed system. Please note that we are not permitted to disclose the image data used in this experiment. Each image was an 8-bit grayscale with a resolution of 2,048 $\times$ 2,048 pixels. Based on expert annotations of the gastric cancer locations, the smallest rectangle surrounding each location was defined as a bounding box to be detected.
A subject-based five-fold cross-validation was used for training and evaluation, with precision, recall, and F1-score serving as the performance indicators for this system. The system provided candidate lesion bounding boxes, with their associated confidence scores exceeding a predefined detection threshold $\xi$. In the subject-based five-fold cross-validation, subjects were randomly divided into five groups. Images from four groups were used for training, while images from the remaining group were used for evaluation. This process was repeated five times, with a different group selected for evaluation each time. In medical research, including this experiment, data from the same subject can be inherently similar. Using standard cross-validation could lead to a virtual data leak, resulting in an overestimation of model performance due to the similarity between training and test data. To prevent this and ensure a rigorous evaluation, we used this evaluation strategy.
@@ 4.2 Details of the detection model
We used EfficientDet-D7 (Tan et al., 2020) as a gastric cancer detection model with an input resolution of 1,536 $\times$ 1,536 pixels. One difference from the original image size was the margin for random cropping performed during training. The entire network was trained end-to-end, allowing it to simultaneously estimate the detection target at any location and size within the input image and classify its content (i.e., ’cancer’). For details on the model, please refer to the original paper. The learning rate was set to 0.001, with a batch size of 32 and 100 training epochs. The optimization method used was momentum SGD (Qian, 1999), with the exponential decay rate of the first-order moment set to 0.9. To construct the proposed gastric cancer detection system, we applied the proposed R-sGAIA as an online augmentation technique during training. Additionally, healthy control images without cancer lesions, which are typically not used in general object detection models, were incorporated into the training process using HBBT.
@@ 4.3.1 Determination of hyperparameters in R-sGAIA
In this section, we describe the hyperparameters $\theta$ and $\gamma$ of the proposed R-sGAIA. $\theta$ is a hyperparameter that determines the edge strength $E(x,y)$ for which the probability of being considered a gastric fold is 50% in the sigmoid function $p(e=E(x,y))$ and was determined to be 0.55 based on the distribution of validation results. For $\gamma$, which characterizes the slope of $p(e)$, we selected $\gamma=4$ from 2, 4, 6, 8, and 10 (Table 1) from the detection performance on validation images.
@@
$\gamma$ Precision (%) Recall (%) F1-score
@@ 4.3.2 Detection of suspicious cancerous regions
To confirm the effectiveness of the proposed system with R-sGAIA and HBBT, we conducted comparative experiments using the following combinations of conditions. Since there is no directly comparable literature on gastric cancer detection, we implemented Laddha et al.’s (2019) method based on YOLOv3 and compared the performance on our task.
@@
    We selected two deep learning-based object detection models for this task: ResNet-101+Faster R-CNN (model 1) and EfficientDet-D7 (model 2). The first model has been widely used in the literature, including in sGAIA studies. The second model is one of the advanced object detection models frequently employed in many studies, although newer models have since been proposed. The training images used in the experiment to detect candidate cancer regions were grayscale, not color images, and since there was no suitable pre-trained model available, we trained the model from scratch using He’s initial weights He et al. (2015) with the random seed set to ’2021’. RandAugment (Cubuk et al., 2020), one of the state-of-the-art online augmentation methods, was applied to these models and served as the baseline for each (i.e., baseline models 1 and 2, respectively). Color augmentation was omitted because this task involves grayscale images. Based on the detection performance of the validation dataset, 4 of the 13 augment types were selected with an intensity level of 3 (i.e., $M=4$ and $N=3$ according to their notation).
@@
    We compared the performance of each model with and without the proposed techniques. For data augmentation, we used either RandAugment, conventional sGAIA (+sGAIA), or the proposed R-sGAIA (+R-sGAIA). The use of the proposed HBBT was indicated as either applied (+HBBT) or not.
@@ 4.3.3 Computational environment for model training and evaluation
In all experiments, we used a computer equipped with a Xeon E5-2650v4 2.2GHz CPU, 256GB of memory, and a GeForce RTX 2080 Ti GPU, running on Ubuntu 20.04 LTS. PyTorch was used as the deep learning framework. The implementations of Faster R-CNN***https://pytorch.org/vision/master/models/faster_rcnn.html, EfficientDethttps://github.com/rwightman/efficientdet-pytorch, and RandAugmenthttps://pytorch.org/vision/main/generated/torchvision.transforms.RandAugment.html were sourced from the links provided. Please note that the OS versions listed here reflect those in use at the time of the initial submission and are not the latest versions.

### v5

@@ 2.3 Hard negative mining
In recent years, deep learning-based object detection models have also adopted this strategy and reported significant performance improvements by relearning misdetected bounding box regions, including in the medical field (Li et al., 2019; Tang et al., 2019). However, such object detection models typically cannot utilize images that do not contain any detection targets (such as background images or healthy control images) for training because the error function to be minimized is defined based on the position and size of the bounding box for each detection target. Existing hard sample mining methods have been applied to data with at least one annotation box of the target class and have used overdetected or undetected regions or image patches as hard samples. Our HBBT is a form of hard sample mining or hard negative mining that retrains regions of false positives. However, it differs significantly in that it allows the use of healthy control cases for training, which could not be used in traditional object detection models. This is a particularly important contribution in the medical field, where the cost of acquiring positive samples is high.
3 Practical gastric cancer diagnostic support system
The system highlights suspected cancerous sites with bounding boxes, thus allowing physicians to confirm the validity of the estimations. The two technical innovations in this study significantly improved system performance and achieved practical accuracy. The overall structure of the proposed system is shown in Figure 1. This study was approved by the Institutional Review Board of Tokai University Hospital (No. 20R-033, approval date: June 12, 2020). The main technical contributions were (1) R-sGAIA, which generates a diverse set of images that effectively emphasize the gastric fold region for data augmentation, and (2) HBBT, which reduces false positives by assigning “hard sample” classes to bounding box regions where false positives are detected and iteratively learning from them. The object detection model used is EfficientDet (Tan et al., 2020), which incorporates EfficientNet (Tan and Le, 2019) as its backbone, a CNN known for its excellent performance, although the choice of model was arbitrary. The input images were 2,048 $\times$ 2,048 pixels, and the output consisted of a bounding box surrounding the presumed cancerous region in the image along with its confidence score.
Figure 1: Overview of the proposed gastric cancer diagnostic support system.
@@ 3.1 R-sGAIA
The proposed R-sGAIA is a refined version of our previous method, sGAIA (Okamoto et al., 2019). Both R-sGAIA and sGAIA are online data augmentation techniques developed based on medical knowledge to generate diverse images that emphasize gastric folds in X-ray images, aiming to detect inflammation on the gastric mucosal surface. The original sGAIA improved gastric cancer detection by 6.9% in terms of the F1 score (recall 92.3%, precision 32.4%) when applied to a Faster R-CNN model with ResNet-101 as the backbone, thereby validating its effectiveness for practical decision support. However, in sGAIA, the emphasis probability of each pixel $p(x,y)$, crucial for effective enhancement, was subjectively determined as a discrete value based on preliminary experiments. In the proposed R-sGAIA, $p(x,y)$ is instead modeled as a flexible continuous function determined by only two hyperparameters. A schematic of the proposed R-sGAIA process is shown in Figure 2. Similar to sGAIA, R-sGAIA consists of four steps, with step 2 being the only difference.
@@
    Step 1: Calculation of edge strength $E(x,y)$
    First, a contrast-enhanced image $I_{e}(x,y)$ is generated from the original gastric X-ray image $I(x,y)$ using histogram equalization. Then, its gradient and high-frequency components, $\nabla I_{e}(x,y)$, and $H_{e}(x,y)$, are obtained using general image processing techniques, such as Canny edge detection filter and Butterworth high-pass filter, respectively. Each component is normalized to [0, 1], and the normalized edge strength $E(x,y)$ is calculated as follows:
@@
    Step 2: Calculation of the probability of a gastric fold region $p(x,y)$
@@
    The parameters $\gamma$ and $\theta$ are hyperparameters that define the slope of the sigmoid function and adjust the probability of running the augmentation of $E(x,y)$ to be 50%. They are determined based on the diagnostic performance of the validation data, as described below. Figure 3 compares the enhancement selection probability $p(e=E(x,y))$ based on the feature intensity $e$ for R-sGAIA and sGAIA. Note that, in sGAIA, the probability $p(e)$ is subjectively determined based on preliminary experiments, as described above.
@@
    Step 3: Determination of gastric fold edge region $G(x,y)$
@@
    Step 4: Generation of enhanced gastric fold images $A(x,y)$
@@
Here, $\alpha$ and $\beta$ are augmentation parameters. The range of $\alpha$ was from 0.9 to 1.0, and $\beta$ was from --15 to --5, as determined by prior experiments. During runtime, these parameters are randomly selected from within their respective ranges. Using these parameters, the proposed R-sGAIA probabilistically enhances the intensity of the gastric fold region each time to generate a different enhanced gastric image for each trial. This makes it highly compatible with online data augmentation in deep neural network training. Another significant advantage of R-sGAIA is that it does not require any additional runtime. We applied this R-sGAIA for online augmentation to train a machine learning model (EfficientDet-D7) used for cancer region detection.
@@
Figure 4: Positive example and five typical false positives provided by experts (as shown in red box with “malignant” label. In HBBT, these false detection regions are labeled as the “hard boundary box” class, and the model is retrained accordingly.).
@@ 3.2 HBBT
The task of detecting cancerous regions in gastric X-ray images is a one-class detection problem that focuses exclusively on positive (i.e., malignant) regions. However, due to the presence of numerous regions with similar visual features, distinguishing these regions is challenging not only for machine learning models but also for experts. Figure 4 illustrates typical false detections identified in experiments using the EfficientDet-D7 object detection model. The results indicate that most of these false detections can be categorized into five categories, as shown in examples (b)–(f). However, it is difficult to predefine and control these features explicitly. Reducing such hard negatives–non-target regions with high visual similarity–is crucial. As previously mentioned, existing deep learning-based object detection algorithms generally lack the capability to explicitly learn from readily available negative data, which presents a significant limitation, particularly in medical tasks such as this.
The concept of HBBT is illustrated on the right side of Figure 1. HBBT extends the concept of hard negative mining by more effectively identifying regions prone to false negatives. It achieves this by incorporating healthy data that could not previously be utilized for training and retraining the model to avoid detecting these regions. In essence, HBBT leverages healthy data that was previously unavailable for training to reduce false negatives in regions where the model might otherwise misclassify with high confidence, thereby improving detection accuracy. The concrete procedure of HBBT for any object detection model $M$ is as follows:
  • 1.
    Step 1: Train the model $M$ using the original labeled training data where the ground truth bounding boxes correspond to the target malignant label.
  • 2.
    Step 2: Perform the box detection procedures using model $M$ on healthy (non-target) data. Each falsely detected bounding box region (e.g., as shown in Figure 4, (b)–(f)) is labeled as a “hard sample” class. We refer to these false detection boxes as hard boundary boxes, and the images containing them are subsequently added to the training data.
  • 3.
    Step 3: The model $M$ is re-trained with updated training data for the two-class detection task of “malignant” and “hard samples.”
Note that the “hard sample” class represents a non-cancerous class and is ignored during the execution of model $M$, thus treating its output as negative.
@@ 4.1 Datasets and evaluation criteria
The image set used in this study consisted of 4,724 gastric X-ray images (1,117 images with 1,504 lesion annotations by radiologists and 3,607 images without lesions) from 145 patients in clinical settings provided by the Tokai University School of Medicine, Japan. A total of 116 images from five randomly selected patients (49 of them with 49 annotated boxes) were used as validation data to determine the hyperparameters and to decide when to terminate HBBT. The remaining 4,608 images from 140 patients were used to train and evaluate the proposed system. Please note that we are not permitted to disclose the image data used in this experiment. Each image was an 8-bit grayscale with a resolution of 2,048 $\times$ 2,048 pixels. Based on expert annotations of the gastric cancer locations, the smallest rectangle surrounding each location was defined as a bounding box to be detected.
A subject-based five-fold cross-validation was used for training and evaluation, with precision, recall, and F1 score serving as the performance indicators for this system. The system provided candidate lesion bounding boxes with associated confidence scores exceeding a predefined detection threshold $\xi$. In the subject-based five-fold cross-validation, subjects were randomly divided into five groups. Images from four groups were used for training, while images from the remaining group were used for evaluation. This process was repeated five times, with a different group selected for evaluation each time. In medical research, including this experiment, data from the same subject can be inherently similar. Using standard cross-validation could lead to a virtual data leak and result in an overestimation of model performance due to the similarity between training and test data. To mitigate this and ensure rigorous evaluation, we adopted the subject-based evaluation strategy.
@@ 4.2 Details of the detection model
We used EfficientDet-D7 (Tan et al., 2020) as a gastric cancer detection model with an input resolution of 1,536 $\times$ 1,536 pixels. One difference from the original image size was the margin for random cropping performed during training. The entire network was trained end-to-end, which allowed it to simultaneously estimate the detection target at any location and size within the input image and classify its content (i.e., “cancer”). For details on the model, please refer to the original paper. The learning rate was set to 0.001, with a batch size of 32 and 100 training epochs. The optimization method used was momentum stochastic gradient descent (SGD) (Qian, 1999), with the exponential decay rate of the first-order moment set to 0.9. To construct the proposed gastric cancer detection system, we applied the proposed R-sGAIA as an online augmentation technique during training. Additionally, healthy control images without cancer lesions, which are typically not used in general object detection models, were incorporated into the training process using HBBT.
@@ 4.3.1 Determination of hyperparameters in R-sGAIA
In this section, we describe the hyperparameters $\theta$ and $\gamma$ of the proposed R-sGAIA. $\theta$ is a hyperparameter that determines the edge strength $E(x,y)$ for which the probability of being considered a gastric fold is 50% in the sigmoid function $p(e=E(x,y))$ and was determined to be 0.55 based on the distribution of the validation results. For $\gamma$, which characterizes the slope of $p(e)$, we selected $\gamma=4$ from 2, 4, 6, 8, and 10 (Table 1) from the detection performance on the validation images.
@@
$\gamma$ Precision (%) Recall (%) F1 score
@@
4.3.2 Determination of hyperparameter in HBBT
The detection threshold for hard boundary boxes in HBBT is set to match the detection threshold for positive boundary boxes across the entire model $\xi$. In this study, the parameter $\xi$ was determined to achieve a sensitivity of approximately 90% for detecting candidate cancer regions, as detailed below. The primary goal of HBBT is to reduce false positives that the model detects with high confidence. When applying HBBT, a higher detection threshold generates high-quality hard boundary boxes. However, if the threshold is set too high, the number of hard boundary boxes becomes limited, reducing the overall effectiveness of HBBT. Conversely, if the threshold is set too low, it may produce an excessive number of hard boundary boxes unrelated to false positives, potentially interfering with the detection of critical cancer regions near these boxes. In preliminary experiments where $\xi$ was set to this value, the best results were achieved when the threshold for detecting hard boundary boxes was equal to the threshold for detecting cancer regions, $\xi$.
@@ 4.3.3 Detection of suspicious cancerous regions
To confirm the effectiveness of the proposed system with R-sGAIA and HBBT, we conducted comparative experiments using the following combinations of conditions. Because there is no directly comparable literature on gastric cancer detection, we implemented Laddha et al.’s (2019) method based on YOLOv3 and compared the performance on our task.
@@
    We selected two deep learning-based object detection models for this task: ResNet-101 + Faster R-CNN (model 1) and EfficientDet-D7 (model 2). The first model has been widely used in the literature, including in sGAIA studies. The second model is one of the advanced object detection models frequently employed in many studies, although newer models have since been proposed. The training images used in the experiment to detect candidate cancer regions were grayscale, not color images. Also, because there was no suitable pre-trained model available, we trained the model from scratch using He’s initial weights (He et al., 2015) with the random seed set to “2021”. RandAugment (Cubuk et al., 2020), one of the state-of-the-art online augmentation methods, was applied to these models and served as the baseline for each (i.e., baseline models 1 and 2). Color augmentation was omitted because this task involved grayscale images. Based on the detection performance of the validation dataset, 4 of the 13 augment types were selected with an intensity level of 3 (i.e., $M=4$ and $N=3$ as defined in the original paper). The configuration of these machine learning models and training techniques is identical to that presented in the original paper.
@@
    We compared the performance of each model with and without the proposed techniques. For data augmentation, we used either RandAugment, conventional sGAIA (+sGAIA), or the proposed R-sGAIA (+R-sGAIA). The use of the proposed HBBT was indicated as either applied (+HBBT) or not applied.
@@ 4.3.4 Computational environment for model training and evaluation
In all of the experiments, we used a computer equipped with a Xeon E5-2650v4 2.2GHz CPU, 256GB of memory, and a GeForce RTX 2080 Ti GPU, running on Ubuntu 20.04 LTS. PyTorch was used as the deep learning framework. The implementations of Faster R-CNN***https://pytorch.org/vision/master/models/faster_rcnn.html, EfficientDethttps://github.com/rwightman/efficientdet-pytorch, and RandAugmenthttps://pytorch.org/vision/main/generated/torchvision.transforms.RandAugment.html were sourced from the links provided. Please note that the OS versions listed here reflect those in use at the time of the initial submission and were not the latest versions.